OpenStudy (anonymous):

What is a slope of a line that passes through points ( -2,3) and ( 4,2)

OpenStudy (anonymous):

the formula for finding the slope is (y2 - y1)/(x2-x1).

OpenStudy (anonymous):

4+2/2-3=6/(-1)=-6

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

oops sorry, its -1/6

OpenStudy (anonymous):

Can you show mehow you got that